# ExteriorExtensions: A package for Macaulay2

### v1

@@ 2. Examples
The algebras $\mathfrak{a}=\mathfrak{sl}_{2m}\oplus{\textstyle\bigwedge^{m}}\mathbb{C}^{2m}$ for $m=2$ and $m=4$ are actually simple Lie algebras, respectively $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}(6)$ and $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$. An excellent overview on fine gradings of simple Lie algebras is given in [MR3601079]. It is not immediately clear to the author if this example falls within those fine gradings classified in [MR3601079, Example 6].
@@
We have a sporadic isomorphism $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}_{6}$. Recall from [FultonHarris, Ch. 15.2 and Ch.17.1] the structure of the two algebras. The weights of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ are the vertices of an octahedron, while the weights of $\mathfrak{sl}_{4}$ are the midpoints of the edges of a cube. These geometries are also present in the weights of a basis of $\mathfrak{sp}_{6}$, see [FultonHarris, p220, p254]. We set the algebra as an ExteriorExtension via:
i1 : loadPackage"ExteriorExtensions"
o1 = ExteriorExtensions
o1 : Package
i2 : ea24 = buildAlgebra(2,4,e,QQ)
o2 = ea24
o2 : ExteriorExtension
We will see that $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ with the (essentially unique) bracket coming from the GJD criterion forms a Lie algebra. We can check (at least on random points) if the bracket satisfies the Lie algebra axioms in each graded piece of the algebra. For grade 0 times grade 0 we perform the following:
i3 : A = makeTraceless random(QQ^4,QQ^4)
o3 = | 443/168 9/2    3/2      3/5    |
     | 4       23/168 4        3/7    |
     | 3/2     5/2    -265/168 9/5    |
     | 9       1/3    5/8      -67/56 |
              4        4
o3 : Matrix QQ  <--- QQ
i4 : B = makeTraceless random(QQ^4,QQ^4)
o4 = | 13/280 1/4     3/5     2/3     |
     | 2/3    -71/280 1/10    3/10    |
     | 1/5    5/2     -51/280 5/3     |
     | 1      2/5     3/5     109/280 |
              4        4
o4 : Matrix QQ  <--- QQ
i5 : ea24.bracket(A,B) + ea24.bracket(B,A)
o5 = 0
              4        4
o5 : Matrix QQ  <--- QQ
i6 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o6 = 0
              21        21
o6 : Matrix QQ   <--- QQ
@@
For a pair of grade 1 elements we perform the following:
i7 : A = random(2, ea24.appendage)
     4       5       7               5
o7 = -e e  + -e e  + -e e  + 5e e  + -e e  + 2e e
     7 0 1   9 0 2   9 1 2     0 3   6 1 3     2 3
o7 : QQ[e ..e ]
         0   3
i8 : B = random(2, ea24.appendage)
     1       3       2       2       1       10
o8 = -e e  + -e e  + -e e  + -e e  + -e e  + --e e
     5 0 1   2 0 2   3 1 2   9 0 3   3 1 3    3 2 3
o8 : QQ[e ..e ]
         0   3
i9 : ea24.bracket(A,B) + ea24.bracket(B,A)
o9 = 0
                        4                  4
o9 : Matrix (QQ[e ..e ])  <--- (QQ[e ..e ])
                 0   3              0   3
i10 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o10 = 0
               21        21
o10 : Matrix QQ   <--- QQ
For a point in grade 0 and a point in grade 1 we see the axioms hold, but we recall the caveat that it is important to make the matrix traceless, otherwise the Jacobi identity fails.
i11 : A = makeTraceless random(QQ^4,QQ^4)
o11 = | 17/80 10/9  7/2    1/2  |
      | 5/9   21/80 3/7    9/5  |
      | 1     4/5   -43/80 3/2  |
      | 1/4   3/2   3      1/16 |
               4        4
o11 : Matrix QQ  <--- QQ
i12 : B = random(2, ea24.appendage)
             9       2              8       10
o12 = e e  + -e e  + -e e  + e e  + -e e  + --e e
       0 1   5 0 2   9 1 2    0 3   9 1 3    9 2 3
o12 : QQ[e ..e ]
          0   3
i13 : ea24.bracket(A,B) + ea24.bracket(B,A)
o13 = 0
o13 : QQ[e ..e ]
          0   3
i14 : ea24.ad(ea24.bracket(A,B)) - ea24.bracket(ea24.ad(A),ea24.ad(B))
o14 = 0
               21        21
o14 : Matrix QQ   <--- QQ
We can also construct a Killing matrix and see that it is $21\times 21$ and non-degenerate. The only possibility for a simple Lie algebra over $\mathbb{C}$ of dimension 21 is $\mathfrak{sp}_{6}$. One checks that the Killing matrix has a full set of real eigenvectors, so the isomorphism $\mathfrak{a}\cong\mathfrak{sp}_{6}$ holds over $\mathbb{R}$.
i15 : K = ea24.KillingMatrix()
o15 = | 20 10 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 10 20 10 0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
      | 0  10 20 0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   0  |
@@
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  0   20 |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0  -20 0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   0  20 0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0   20 0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  -20 0  0  0   0  |
      | 0  0  0  0  0  0  0  0  0  0  0  0  0  0  0  20 0   0  0  0   0  |
               21        21
o15 : Matrix QQ   <--- QQ
i16 : rank K
o16 = 21
i17 :  toList eigenvalues K
o17 =  {34.1421, 20, 5.85786, 10, -10, 10, -10, 10, -10, 10,
          -10, 10, -10, 10, -10, 20, -20, 20, -20, 20, -20}
o17 : List
The tensors $e_{i}\wedge e_{j}+e_{k}\wedge e_{l}$ for $i,j,k,l$ distinct have diagonalizable adjoint forms (we call them semi-simple or ad-semi-simple). We can check that the geometric and algebraic multiplicities of the eigenvalues add up:
i18 : A = ea24.ad(e_0*e_1 + e_2*e_3)
o18 = | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 -1 0  0  0 0  1  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 2  0  0  0 0  -2 |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 -1 0  0  0 0  1  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  -2 0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  2  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 -2 0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  2 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  2 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 2  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  -2 0  0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  -2 0 0  0  |
      | 0 0  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 1  0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  1 0  0 0 0 0 0 -1 0  0 0  0  0  0 0  0  |
      | 0 0  0 0 -1 0 0  0 0 0 0 0 0  -1 0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 0  1 0 0 1 0 0  0  0 0  0  0  0 0  0  |
      | 0 0  0 0 0  0 -1 0 0 0 0 1 0  0  0 0  0  0  0 0  0  |
      | 0 -1 0 0 0  0 0  0 0 0 0 0 0  0  0 0  0  0  0 0  0  |
              21        21
o18 : Matrix QQ   <--- QQ
i19 : round\ toList eigenvalues A
o19 = {0, 0, 2, -2, -2, 0, 2, -2, 0, -2, 2, 2, -2, 2, 0, 0, 0, 0, 0, 0, 0}
o19 : List
i20 : rank A
o20 = 10
The semi-simple structure is also indicated by the block ranks computation:
i21 :  ea24.prettyBlockRanks A
o21 = |g00|g01|g10|g11|total|
@@
The rows in the above table correspond to subsequent powers of the matrix.
@@
i22 : ea24.bracket(e_0*e_1 + e_2*e_3,e_0*e_2 + e_1*e_3)
o22 = {1} | 0 0  0  2 |
      {1} | 0 0  -2 0 |
      {1} | 0 -2 0  0 |
      {1} | 2 0  0  0 |
                         4                  4
o22 : Matrix (QQ[e ..e ])  <--- (QQ[e ..e ])
                  0   3              0   3
i23 : B = ea24.ad(e_0*e_2 + e_1*e_3);
               21        21
o23 : Matrix QQ   <--- QQ
i24 : 0*ea24.bracket(A,B)== ea24.bracket(A,B)
o24 = false
In fact, we checked that the only elements of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ that commute with $e_{i}\wedge e_{j}+e_{k}\wedge e_{l}$ are scalar multiples of itself. So we can’t hope for a Cartan subalgebra that intersects ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ and includes more than one of these forms.
i25 : L = apply(ea24.bases#1, xx-> sub(ea24.bracket(xx,e_0*e_1 + e_2*e_3), QQ))
o25 = {| 1 0 0  0  |, | 0 0 0 -2 |, | 0 0 2 0 |, | 0  0 0 0  |, | 0  0 0 0 |, | -1 0  0 0 |}
       | 0 1 0  0  |  | 0 0 0 0  |  | 0 0 0 0 |  | 0  0 0 -2 |  | 0  0 2 0 |  | 0  -1 0 0 |
       | 0 0 -1 0  |  | 0 2 0 0  |  | 0 0 0 0 |  | -2 0 0 0  |  | 0  0 0 0 |  | 0  0  1 0 |
       | 0 0 0  -1 |  | 0 0 0 0  |  | 0 2 0 0 |  | 0  0 0 0  |  | -2 0 0 0 |  | 0  0  0 1 |
o25 : List
i26 : R = QQ[x_0.. x_5]
o26 = R
o26 : PolynomialRing
i27 : decompose ideal flatten entries sum(6, i-> sub(L_i, R)*x_i)
o27 = {ideal (x , x , x , x , x  - x )}
@@
o27 : List
i28 : first ea24.bases#1 + last ea24.bases#1
o28 = e e  + e e
       0 1    2 3
o28 : QQ[e ..e ]
          0   3
@@
We consider $\mathfrak{sl}_{6}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{6}$, which is not a Lie algebra as it fails to be skew-commuting and it fails the Jacobi identity.
i2 : ea36 = buildAlgebra(3,6,QQ);
i3 : A = random(3, ea36.appendage);
i4 : B = random(3, ea36.appendage);
i5 : ea36.bracket(A,B) + ea36.bracket(B,A) == 0
o5 = false
i6 : ea36.bracket(A,B) - ea36.bracket(B,A) == 0
o6 = true
i7 : ea36.ad(ea36.bracket(A,B)) - ea36.bracket(ea36.ad(A),ea36.ad(B)) == 0
o7 = false
@@
Consider the case of $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$ considered in [Antonyan, oeding2022]. Antonyan shows that a 7-dimensional Cartan subalgebra is given by 4-forms of type $e_{I}+e_{J}$ with the concatenation $I|J$ a permutation of $[8]$, the 7 pairs of indices $I,J$ as follows:
$\begin{matrix}\{\{1,2,3,4\},\;\{5,6,7,8\}\},&&\{\{1,3,5,7\},\;\{6,8,2,4\}\},\\ \{\{1,5,6,2\},\;\{8,4,7,3\}\},&&\{\{1,6,8,3\},\;\{4,7,5,2\}\},\\ \{\{1,8,4,5\},\;\{7,2,6,3\}\},&&\{\{1,4,7,6\},\;\{2,3,8,5\}\},\\ \{\{1,7,2,8\},\;\{3,5,4,6\}\}\end{matrix}$
We can check the diagonalizability and commutativity as follows.
i2 : ea48 = buildAlgebra(4,8,e,QQ);
i3 : list2E = L-> sum(L, l-> product(l, i-> e_(i-1)));
i4 : aa = {{{1,2,3,4},{5,6,7,8}}, {{1,3,5,7},{6,8,2,4}},{{1,5,6,2},{8,4,7,3}},
@@
i5 : for i to length(aa)-1 do(
@@
i6 : matrix apply(7, i-> apply(7, j-> rank ea48.bracket(adList_i,adList_j) ))
o6 = 0
              7        7
o6 : Matrix ZZ  <--- ZZ
@@
Larger algebras are also possible, such as $\mathfrak{e}_{8}\cong\mathfrak{sl}_{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{9}\oplus{\textstyle\bigwedge^{6}}\mathbb{C}^{9}$ from [Vinberg-Elashvili], and we have done extensive computations in [HolweckOeding22]. Here are some example computations that check that the bracket satisfies the Lie-bracket conditions (skew-symmetry and Jacobi).
i2 :     ea = buildAlgebra(3,9)
o2 = ea
o2 : ExteriorExtension
i3 :     E = ea.appendage
o3 = E
o3 : PolynomialRing, 9 skew commutative variables
i4 :     A = makeTraceless random(QQ^9,QQ^9);
              9        9
o4 : Matrix QQ  <--- QQ
i5 :     B = makeTraceless random(QQ^9,QQ^9);
              9        9
o5 : Matrix QQ  <--- QQ
i6 :     ea.bracket(A,B) + ea.bracket(B,A)
o6 = 0
              9        9
o6 : Matrix QQ  <--- QQ
i7 :     ea.ad(ea.bracket(A,B)) -ea.bracket(ea.ad(A),ea.ad(B))
   A = makeTraceless random(QQ^9,QQ^9);
o7 = 0
              248        248
o7 : Matrix QQ    <--- QQ
i8 :
              9        9
o8 : Matrix QQ  <--- QQ
i9 :    B = random(3, E);
i10 :    ea.bracket(A,B) + ea.bracket(B,A)
o10 = 0
o10 : E
i11 :    t1 =    ea.ad(ea.bracket(A,B));
               248        248
o11 : Matrix QQ    <--- QQ
i12 :    t2 =  ea.bracket(ea.ad(A),ea.ad(B));
               248        248
o12 : Matrix QQ    <--- QQ
i13 :    t1-t2   -- gives zero!
o13 = 0
               248        248
o13 : Matrix QQ    <--- QQ
i14 :    A = makeTraceless random(QQ^9,QQ^9);
               9        9
o14 : Matrix QQ  <--- QQ
i15 :    B = random(6, E);
i16 :    ea.bracket(A,B) + ea.bracket(B,A)
o16 = 0
o16 : E
i17 :    ea.ad(ea.bracket(A,B)) - ea.bracket(ea.ad(A),ea.ad(B)) -- gives zero!
o17 = 0
               248        248
o17 : Matrix QQ    <--- QQ
i18 :    A = random(3, E);
i19 :    B = random(6, E);
i20 :    ea.bracket(A,B) + ea.bracket(B,A)
o20 = 0
              9       9
o20 : Matrix E  <--- E
i21 :    t1 =    ea.ad(ea.bracket(A,B));
               248        248
o21 : Matrix QQ    <--- QQ
i22 :    t2 =  ea.bracket(ea.ad(A),ea.ad(B));
               248        248
o22 : Matrix QQ    <--- QQ
i23 :    rank(t1-t2) -- gives zero!
o23 = 0
i24 :    A = makeTraceless random(QQ^9,QQ^9);
               9        9
o24 : Matrix QQ  <--- QQ
i25 :    pt = random(3, E);
During the development of this code we found a problem with our definition of the bracket that stemmed back to our implementation of the Hodge star operator, which we were able to fix. The following equations 2.5 and 2.6 from [Vinberg-Elashvili]
i26 :    ea.bracket(A, ea.star(pt )) + ea.star ea.bracket(transpose A, pt)
o26 = 0
o26 : E
i27 :    pt = E_0*E_1*E_2; tmp  = ea.bracket(pt, ea.star(pt))
o28 = {2} | -2/3 0    0    0   0   0   0   0   0   |
@@
o29 : E

### v2

@@ 2. Examples
The algebras $\mathfrak{a}=\mathfrak{sl}_{2m}\oplus{\textstyle\bigwedge^{m}}\mathbb{C}^{2m}$ for $m=2$ and $m=4$ are actually simple Lie algebras, respectively $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}_{6}$ (a symplectic Lie algebra) and $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$ (an exceptional Lie algebra). An excellent overview of fine gradings of simple Lie algebras is given in [MR3601079]. It is not immediately clear to the author if this example falls within those fine gradings classified in [MR3601079, Example 6].
@@
In this example we show the use of the functions following, buildAlgebra, makeTraceless, bracket, ad, appendage, and KillingMatrix, to illustrate the known sporadic isomorphism $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\cong\mathfrak{sp}_{6}$. Recall from [FultonHarris, Ch. 15.2 and Ch.17.1] the structure of the two algebras. The weights of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ are the vertices of an octahedron, while the weights of $\mathfrak{sl}_{4}$ are the midpoints of the edges of a cube. These geometries are also present in the weights of a basis of $\mathfrak{sp}_{6}$, see [FultonHarris, p220, p254].
In this example we construct $\mathfrak{a}=\mathfrak{sl}_{4}(\mathbb{F})\oplus{\textstyle\bigwedge^{2}}\mathbb{F}^{4}$ for $\mathbb{F}=\mathbb{Q}$, and note some of the salient features of this algebra extension, even though we may be mainly interested in the theory of the complex Lie algebras. Near the end of the example, we note that the Killing form of the algebra has full rank and all real eigenvalues, and this means that the isomorphism $\mathfrak{a}\cong\mathfrak{sp}_{6}$ holds as in isomorphism of real Lie algebras.
We set the algebra as an ExteriorExtension via:
     loadPackage"ExteriorExtensions"
     ea24 = buildAlgebra(2,4,e,QQ)
We will see that $\mathfrak{sl}_{4}\oplus{\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ with the (essentially unique) bracket coming from the GJD criterion forms a Lie algebra. We can check (at least on random points with rational coefficients) if the bracket satisfies the Lie algebra axioms in each graded piece of the algebra. If we had the patience we could also check these identities for all pairs of basis vectors. For illustration we only perform tests for randomly chosen elements. For grade 0 times grade 0, we first make 2 random traceless matrices:
     A0 = makeTraceless random(QQ^4,QQ^4)
     B0 = makeTraceless random(QQ^4,QQ^4)
To access the bracket and adjoint operator ad for the algebra ea24 that we have constructed, we do so with the dot, which we see in the following random check of the skew symmetry and Jacobi identities:
     ea24.bracket(A0,B0) + ea24.bracket(B0,A0)
          0
     ea24.ad(ea24.bracket(A0,B0)) - ea24.bracket(ea24.ad(A0),ea24.ad(B0))
          0
@@
For a pair of grade 1 elements we call for random elements of the appendage and check their brackets are skew and satisfy the Jacobi identity as follows:
     A1 = random(2, ea24.appendage)
     B1 = random(2, ea24.appendage)
     ea24.bracket(A1,B1) + ea24.bracket(B1,A1)
          0
     ea24.ad(ea24.bracket(A1,B1)) - ea24.bracket(ea24.ad(A1),ea24.ad(B1))
          0
For a point in grade 0 and a point in grade 1, we see the axioms hold, but we recall the caveat that it is important to make the matrix traceless (we implemented this projection as makeTraceless), otherwise the Jacobi identity fails. We proceed as before:
     ea24.bracket(A0,B1) + ea24.bracket(B1,A0)
          0
     ea24.ad(ea24.bracket(A0,B1)) - ea24.bracket(ea24.ad(A0),ea24.ad(B1))
          0
We can also construct a Killing matrix and see that it is $21\times 21$ and non-degenerate. The only possibility for a simple Lie algebra over $\mathbb{C}$ of dimension 21 is $\mathfrak{sp}_{6}$.
     K = ea24.KillingMatrix()
          | 20  -10 0   0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   0   0  0   |
          | -10 20  -10 0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   0   0  0   |
          | 0   -10 20  0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   0   0  0   |
@@
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   0   0  -10 |
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   0   10 0   |
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  0   0  0   -10 0  0   |
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  0   0  -10 0   0  0   |
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  0   10 0   0   0  0   |
          | 0   0   0   0  0  0  0  0  0  0  0  0  0  0  0  -10 0  0   0   0  0   |
     rank K
          21
We implemented the KillingMatrix as a function rather than something that gets constructed when the algebra is constructed since, without additional tricks, it may be time-consuming to construct this matrix. One checks that the Killing matrix has a full set of real eigenvectors, so the isomorphism $\mathfrak{a}\cong\mathfrak{sp}_{6}$ holds over $\mathbb{R}$.
     toList eigenvalues K
          {34.1421, 20, 5.85786, 10, -10, 10, -10, 10, -10, 10, -10, 10,
           -10, 10, -10, 10, -10, 10, -10, 10, -10}
Here we note the first instance of a change in the field of definition. To compute eigenvalues M2 automatically upgraded the matrix $K$, or at least its eigenvalues, to work over $\mathbb{C}$.
The tensors $e_{i}\wedge e_{j}\pm e_{k}\wedge e_{l}$ for $i,j,k,l$ distinct have diagonalizable adjoint forms (we call them semi-simple or ad-semi-simple). We can check that the geometric and algebraic multiplicities of the eigenvalues add up:
     E = ea24.appendage
     A = ea24.ad(E_0*E_1 - E_2*E_3)
     tally apply(toList eigenvalues (A), xx-> round(2, realPart xx) +
     round(2, imaginaryPart xx)*ii)
          Tally{-1.41 => 5}
                0 => 11
                1.41 => 5
     rank A
          10
We also note that even though $A$ is defined over QQ, we can let M2 work with square roots.
     tally apply(toList eigenvalues (1/sqrt(2)*A), xx->
     round(2, realPart xx) + round(2, imaginaryPart xx)*ii)
          Tally{-1 => 5}
                0 => 11
                1 => 5
The semi-simple structure is also indicated by the block ranks computation, via blockRanks (to produce a matrix as a result) or via prettyBlockRanks for onscreen printing:
     ea24.prettyBlockRanks A
           |g00|g01|g10|g11|total|
@@
The rows in the above table correspond to subsequent powers of the matrix, and we stop computation after sufficient repetition or zero total rank occurs.
@@
     ea24.bracket(E_0*E_1 + E_2*E_3,E_0*E_2 + E_1*E_3)
          {1} | 0  0 0 -1 |
          {1} | 0  0 1 0  |
          {1} | 0  1 0 0  |
          {1} | -1 0 0 0  |
We checked that the only elements of ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ that commute with $e_{i}\wedge e_{j}+e_{k}\wedge e_{l}$ are scalar multiples of itself. So we can’t hope for a Cartan subalgebra that intersects ${\textstyle\bigwedge^{2}}\mathbb{C}^{4}$ and includes more than one of these forms. We did this computation by accessing the basis (of the grade 1 part) of the algebra via ea24.bases.
     L = apply(ea24.bases#1, xx-> ea24.bracket(xx,e_0*e_1 + e_2*e_3))
Then we prepare additional variables, and solve the system of equations coming from the commutation conditions:
     R = QQ[x_0.. x_5]
     decompose ideal flatten entries sum(6, i-> sub(L_i, R)*x_i)
          ideal (x , x , x , x , x  - x )}
@@
Here we note that we have changed rings again only when necessary. Attention should be paid to the ring of definition for the matrices the package produces, and sometimes substitution commands are necessary. ∎
@@
We consider $\mathfrak{sl}_{6}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{6}$, which is not a Lie algebra as it fails to be skew-commuting (it is commuting on the grade 1 piece) and it fails the Jacobi identity.
     ea36 = buildAlgebra(3,6,QQ);
     A = random(3, ea36.appendage);
     B = random(3, ea36.appendage);
     ea36.bracket(A,B) + ea36.bracket(B,A) == 0
          false
     ea36.bracket(A,B) - ea36.bracket(B,A) == 0
          true
     ea36.ad(ea36.bracket(A,B)) - ea36.bracket(ea36.ad(A),ea36.ad(B)) == 0
          false
Interestingly, the trace of the 4th power of the adjoint operator still produces the single (degree 4) generator of the invariant ring $\mathbb{C}[{\textstyle\bigwedge^{3}}\mathbb{C}^{6}]^{\operatorname{SL}_{6}}\cong\mathbb{C}[f_{4}]$ in this case. Computing values of this invariant is accomplished as follows:
     E = ea36.appendage
     f4 = A-> trace(A^4)
     N = E_0*E_1*E_2 + E_3*E_4*E_5
     f4 ea36.ad(N)
          36
@@
Consider the case of $\mathfrak{sl}_{8}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{8}\cong\mathfrak{e}_{7}$ considered in [Antonyan, oeding2022]. Antonyan shows that a 7-dimensional Cartan subalgebra is given by 4-forms of type $e_{I}+e_{J}$ with the concatenation $I|J$ a permutation of $[8]$, the 7 pairs of indices $I,J$ are the following
$\begin{matrix}\{\{1,2,3,4\},\;\{5,6,7,8\}\},&&\{\{1,3,5,7\},\;\{6,8,2,4\}\},&&\{\{1,5,6,2\},\;\{8,4,7,3\}\},\\ \{\{1,6,8,3\},\;\{4,7,5,2\}\},&&\{\{1,8,4,5\},\;\{7,2,6,3\}\},&&\{\{1,4,7,6\},\;\{2,3,8,5\}\},\\ \{\{1,7,2,8\},\;\{3,5,4,6\}\}.\end{matrix}$
We can check the diagonalizability and commutativity as follows. First, we construct the algebra and our candidate for the basis of the Cartan subalgebra.
     ea48 = buildAlgebra(4,8,e,QQ);
     list2E = L-> sum(L, l-> product(l, i-> e_(i-1)));
     aa = {{{1,2,3,4},{5,6,7,8}},{{1,3,5,7},{6,8,2,4}},{{1,5,6,2},{8,4,7,3}},
@@
Then we produce the adjoint operators for each, and compute respectively their sums of the algebraic and geometric multiplicies of eigenvalues to see that they do, in fact, agree, and hence these elements are diagonalizable.
     for i to length(aa)-1 do(
@@
Next, we check that these elements mutually pairwise commute by computing all their commutators. We store these results as a matrix whose entries are the ranks of the commutators, and note that we get the zero matrix:
     matrix apply(7, i-> apply(7, j-> rank ea48.bracket(adList_i,adList_j)))
          0
@@
Larger algebras are also possible, such as $\mathfrak{e}_{8}\cong\mathfrak{sl}_{9}\oplus{\textstyle\bigwedge^{3}}\mathbb{C}^{9}\oplus{\textstyle\bigwedge^{6}}\mathbb{C}^{9}$ from [Vinberg-Elashvili], and we have done extensive computations in [HolweckOeding23]. Here are some example computations that check that the bracket satisfies the Lie-bracket conditions (skew-symmetry and Jacobi). We check bracket conditions for pairs of random elements in each grade:
     ea = buildAlgebra(3,9)
     E = ea.appendage
     A0 = makeTraceless random(QQ^9,QQ^9);
     B0 = makeTraceless random(QQ^9,QQ^9);
     A1 = random(3, E);
     B1 = random(3, E);
     A2 = random(6, E);
     B2 = random(6, E);
     ea.bracket(A0,B0) + ea.bracket(B0,A0)
          0
     ea.ad(ea.bracket(A0,B0)) -ea.bracket(ea.ad(A0),ea.ad(B0))
          0
     ea.bracket(A0,B1) + ea.bracket(B1,A0)
          0
     ea.ad(ea.bracket(A0,B1)) -ea.bracket(ea.ad(A0),ea.ad(B1))
          0
     ea.bracket(A0,B2) + ea.bracket(B2,A0)
          0
     ea.ad(ea.bracket(A0,B2)) -ea.bracket(ea.ad(A0),ea.ad(B2))
          0
     ea.bracket(A1,B1) + ea.bracket(B1,A1)
          0
     ea.ad(ea.bracket(A1,B1)) -ea.bracket(ea.ad(A1),ea.ad(B1))
          0
     ea.bracket(A1,B2) + ea.bracket(B2,A1)
          0
     ea.ad(ea.bracket(A1,B2)) -ea.bracket(ea.ad(A1),ea.ad(B2))
          0
     ea.bracket(A2,B2) + ea.bracket(B2,A2)
          0
     ea.ad(ea.bracket(A2,B2)) -ea.bracket(ea.ad(A2),ea.ad(B2))
          0
The following match equations 2.5 and 2.6 from [Vinberg-Elashvili].
     pt = E_0*E_1*E_2;
     tmp  = ea.bracket(pt, ea.star(pt))
          {2} | -2/3 0    0    0   0   0   0   0   0   |
@@
3. The Future
What we anticipate for the future of this package is to provide functionality for theoretical work on tensors. For example, it is possible to compute an adjoint form of a tensor of small format in a short amount of time on a laptop computer. From here one may use the adjoint form to compare tensors via their adjoint block ranks, adjoint spectra, etc. This can have applications whenever one wants to separate tensor orbits, for example. This strategy was explored in [HolweckOeding23]. The author has work in progress with collaborators where this strategy is applied to tensors that arise in phylogenetic networks. In particular, we are very interested in the cases beyond the “tame” tensor formats, where there are not finitely many orbits, and the algebra we construct $\mathfrak{a}$ is not a semi-simple Lie algebra. One very interesting case is $\mathfrak{a}=\mathfrak{sl}_{12}\oplus{\textstyle\bigwedge^{4}}\mathbb{C}^{12}\oplus{\textstyle\bigwedge^{8}}\mathbb{C}^{12}$, which contains a graded subalgebra
$\mathfrak{a}_{4,4,4}=(\mathfrak{sl}_{4}\oplus\mathfrak{sl}_{4}\oplus\mathfrak{sl}_{4})\oplus(\mathbb{C}^{4}\otimes\mathbb{C}^{4}\otimes\mathbb{C}^{4})\oplus({\textstyle\bigwedge^{2}}\mathbb{C}^{4}\otimes{\textstyle\bigwedge^{2}}\mathbb{C}^{4}\otimes{\textstyle\bigwedge^{2}}\mathbb{C}^{4})\oplus({\textstyle\bigwedge^{3}}\mathbb{C}^{4}\otimes{\textstyle\bigwedge^{3}}\mathbb{C}^{4}\otimes{\textstyle\bigwedge^{3}}\mathbb{C}^{4}),$
which allows us to utilize these tools to study tensors of format $4\times 4\times 4$. This format is the smallest cubic format for tensors where there is not a complete irredundant classification of orbits. We are quite interested in studying this case further with these new tools.
We have tried to be as efficient as possible in the implementation of the functions in this package, but there is certainly room for improvement, both mathematically and algorithmically. On the mathematical side, if one knows already how to compute the Killing form in a combinatorial fashion it might be possible to skip the lengthy process that computes each entry of the Killing matrix as the trace of a product of adjoint operators. Algorithmically, it might also be possible to speed up the computation of the adjoint operators in the first place, and it also seems very possible to utilize sparse arrays to be more efficient with memory. In addition, it could be quite interesting to attempt to construct similar extensors for other Lie algebras (replacing $\mathfrak{a}_{0}=\mathfrak{sl}_{n}$), or for other modules, replacing $\mathfrak{a}_{<0}$.
4. Acknowledgements
The author thanks Frederic Holweck for his input on the development of this package, which was used for their joint work. Ian Tan also provided useful discussions and testing. The author thanks two anonymous referees for their comments which greatly improved this manuscript. The author acknowledges partial support from the Thomas Jefferson Foundation.